Shift summary for owl, 7/7/99

Jason Hamm (jhamm@physics.arizona.edu)
Wed, 7 Jul 1999 05:02:41 -0700 (MST)


Shift Summary for owl shift, 07/07/99

On Shift: V. Jejer, J. Hamm

Hours of beam available at all intensities: 3.25
Hours of beam available at nominal intensity: 1.75
Hours of E832 data (special runs not included): 3.0
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time: On access, main injector problems

========================================================================
Runs ended this shift:
Run 13832 - BEAM_832 (281K events)
Run 13833 - Laser scan

Tapes brought to FCC this shift: None
========================================================================
Accesses during shift: 1 (see below)
========================================================================
Overview of shift:
Main injector problems. No beam for most of shift - most of what we
get is at low intensity ( <3 E12 )
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
0:00 EB and SET on access to replace repaired HA power
supply

0:11 Onsplit requests QKA002 to be mounted on ktev3.
We happily comply

0:16 End access, start beam832 run 13832

13832 0:30 Beam returns, although MI problems limit SEM to
1 E12, at most

13832 1:10 Beam disappears, no estimate

13832 3:00 Still no beam, still no estimate. Stop run 13832.

3:15 Start laser scan run 13833

3:56 Laser scan finished, still no beam or estimate

4:06 Start beam832 run 13834

13834 4:29 MCR calls and tells us to expect beam soon. 30
seconds later, we start getting low intensity
spills.

13834 4:35 SEM up to 5 E12!!!

13834 4:47 Beam gone again - MI problems. What a tease.

13834 4:54 Beam returns, with low SEM (1-3 E12). We're
warned that it will be intermittent.

13834 5:30 Notice that TW quadrant of ET dropped out. No
audio alarm! Reload ET - quadrant restored.
Lost about 2-5 spills.

Back up to nominal intensity.

13834 6:00 Monitoring histograms show bananas are
inefficient. Resetting doesn't help, Jason L.
en route.

13834 6:45 Jason L. claims inefficiency due to change in TDC
intime window made by DC people last night. Watch
out, DC people - he's on the warpath.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

It might be a good idea to set kdisp to display only laser calibration
triggers (c001), so you can see right away if there's an ET problem. Keep
an eye on the TW quadrant - if it drops out again, reload ET and notify
an expert.